OpenStudy (anonymous):

For which function is f(x) not equal to f^-1(x)? a. f(x)=2-x b. f(x)=2/x c. f(x)=-x d. f(x)=2x please help

OpenStudy (danjs):

to test what f inverse is, swap f(x) and x, and resolve for f(x), that is your f inverse

OpenStudy (danjs):

for example, a. f(x) = 2 - x swap x = 2 - f(x) solve f(x) = 2 - x (same as before) f inverse is the same as f

OpenStudy (danjs):

u get it?

OpenStudy (danjs):

Ill do one more to show you .... b. f(x) = 2/x Swap f(x) and x x = 2/f(x) re-solve for f(x), this is now f inverse of x x*f(x) = 2 f(x) = 2/x f^-1(x) = 2/x F inverse of x is the same as f(x)

OpenStudy (danjs):

try that on c. and d. to see which one changes form when you do that process